# Improving Object Detection by Modifying Synthetic Data with Explainable AI

### v1

@@ 4 Experimental Setup
Our aim is to perform object detection, that is, to detect and classify any vehicle(s) in the image, and estimate the bounding box coordinates of any detected vehicles. We use a YOLOv8 nano (You Only Look Once) model by Ultralytics [53] that has been pretrained on the COCO dataset, due to its small size and high accuracy, and fine-tune it on the ATR DSIAC dataset.
In order to prepare the dataset, we first crop the input images of initial size height 512, width 640 to a size of $256$ by $512$ pixels, with the centre of the crop positioned at a random pixel in the input image, which also provides negative samples where no vehicle is present. The cropped image is then resampled using bilinear interpolation to a size of $128$ by $256$ pixels. We split the images according to vehicle orientations. For our test and real training datasets, we use ATR DSIAC images with vehicle orientations in the ranges $\mathcal{O}_{test}=[70^{\circ},110^{\circ}]\cup[250^{\circ},290^{\circ}]$ and $\mathcal{O}_{train}=[-20^{\circ},20^{\circ}]\cup[160^{\circ},200^{\circ}]$ respectively. For synthetic training data, vehicle orientations are in the range $\mathcal{O}_{syn}=[50^{\circ},130^{\circ}]\cup[230^{\circ},310^{\circ}]$, as illustrated in Fig. 3. We employ this train-test split so that we can examine how the addition and modification of synthetic data improve our model’s ability to detect orientations of the target vehicles that have not been seen in training; detection of unseen scenarios is an application where synthetic data have especial utility. This split reduces data leakage between the train and test data, which can occur if train and test frames are sampled uniformly from the video dataset, as per much of the previous literature on the ATR DSIAC dataset [54, 55, 56, 57, 58]. As neighbouring frames in a video are highly correlated, results reported on such uniformly sampled train-test splits can be unrealistically high due to overfitting.
@@ 5 Results
To provide a baseline reference for improvements in performance we first fine-tune our YOLOv8n model for $120$K iterations using only the real data from the dataset, trained using $9000$ images from $\mathcal{O}_{train}$ and tested using $\mathcal{O}_{test}$, where there is no overlap between the vehicle orientations in $\mathcal{O}_{train}$ and $\mathcal{O}_{test}$. We report performance using the mean average precision score at an intersection over union (IoU) threshold of $0.5$ (mAP50). We take ensemble averages over 4 different random seeds to demonstrate that our model improvements are robust to variations in seed. The baseline performance recorded in the first row of Table 1 shows an mAP50 score of $90.0\%$ when the model is only trained on the real dataset. The second row of this table shows an increase in mAP50 score to $94.6\%$ performance when we instead train our model on the same real data plus 9000 synthetic images generated from our initial mesh models, which we term v0 (version 0). The v0 synthetic data includes orientations ($\mathcal{O}_{syn}$) that overlap the test time orientations $\mathcal{O}_{test}$. Through a suite of tests we show that this increase in performance is not simply due to an increase in dataset size, and that the optimal ratio of real to synthetic data is 0.5 (supplementary Fig. 13), which we adopt through the rest of our experiments.
@@
Real 90.0%
Real + Syn v0 (Initial) 94.6%
Real + Syn vR (Reinforcing) 95.7%
Real + Syn vD (Disruptive) 95.7%
Real + Syn v(R+D) (Reinforcing and Disruptive) 96.1%
@@ 5.1 Using XAI to improve model performance
To understand the increase in mAP50 score from the addition of the v0 synthetic data to the training dataset, we compare the confusion matrices in Fig. 4a,b and see a large reduction in the background being misclassified as vehicles (e.g. dark pink box). By plotting the SHAP values of the pixels, averaged over $100$ random correctly classified test images, we see an increase in focus inside the bounding box of the target vehicle (average SHAP values increase from $39\%$ to $54\%$, see supplementary Fig. 7), which indicates the reduction in misclassification of the background is due to increased model focus on the vehicles relative to the background, when synthetic v0 data are added in training.
However, with the addition of synthetic data v0, the confusion matrices also show a small increase in confusion (decrease in performance) between some vehicles in the dataset e.g. BTR70 misclassified as SUV or ZSU23 (light pink and orange boxes respectively). In the following sections we target these misclassifications using our new procedure from Section 2.1, demonstrating the use of both Reinforcing and Disruptive modifications to the synthetic data v0 in order to reduce model confusion and improve performance.
@@
Figure 4: Confusion matrices calculated by taking an average over $4$ different seeds for models trained with real data plus (a) no synthetic data (b) synthetic data v0, (c) synthetic data vR (reinforcing modification to unique features of SUV relative to BTR70), (d) synthetic data vD (disruptive modification to common features of ZSU23 and BTR70), and (e) synthetic data v(R+D)
@@
Figure 5: Illustration of the synthetic data modification process for reducing confusion. a,d: Ground-truth samples. b,e: Top shows confusion matrix for model trained on the initial synthetic data showing misclassifications highlighted in the confusion matrix. Bottom shows synthetic data samples (left), and SHAP contribution plots (right) for correct classifications capturing the unique features, and for misclassifications that capture the common features. c,f: Top shows confusion matrix for model trained on modified synthetic data, showing reduced misclassifications. Bottom shows samples of modified synthetic data (left), and SHAP contribution plots (right) for correct classifications illustrating an increased focus on unique features and reduced focus on common features. Pixels with a contribution score $<50\%$ (chosen to represent a trade-off between information and noise) of the highest score are masked in purple.
Reinforcing Modification Example: Fig. 5a-c shows how we use Reinforcing Modification to reduce misclassification of the SUV class as a BTR70 (57 instances, light pink box in cropped and full confusion matrix in Fig. 5b and Fig. 4b).
To identify the cause of confusion we group the validation images into $5^{\circ}$ segments of vehicle orientations, which are this dataset’s natural dimension of variation, and plot three average SHAP maps within each cluster; one each for the correct classification of the SUV and BTR70 and one for the misclassification of the SUV as a BTR70.
@@
Correct classifications of the SUV have increased focus towards the rear of the vehicle (Fig. 5b, top green box), whereas misclassifications of the SUV as a BTR70 do not focus in this same area (Fig. 5b, bottom green box) and instead have increased focus on the front bonnet of the SUV, which is a brighter (hotter) area in the IR images. This bright area is in a similar location to the hot rear exhaust of the BTR70 when the SUV is orientated $70^{\circ}-75^{\circ}$ and the BTR70 is orientated $285^{\circ}-290^{\circ}$ (Fig. 5b, red boxes). The bright SUV bonnet and BTR70 exhaust therefore represent common features and the side and rear of the SUV represent unique features. We reinforce model focus on the unique features of the SUV by making the main frame and rear part of the synthetic SUV images more similar to the real SUV images, by reducing the smoothness parameter of the material on the SUV mesh model’s surface. This reduces reflections off the smooth SUV surface (visible as bright-to-dark gradients in Fig. 5b, blue box) and results in more uniform texture (Fig. 5c, blue box), matching the real image (Fig. 5a, top). The effect of this modification is to increase the focus on the unique features and accordingly reduce the focus on common features (Fig. 5c). This results in a reduction of SUV, BTR70 misclassifications from 57 to 24 and an mAP50 increase of 1.1% from 94.6% to 95.7% (Fig. 4b,c, Table 1).
@@
Disruptive Modification Example: Fig. 5d-f shows how we use Disruptive Modification to reduce misclassification of the ZSU23 class as a BTR70 (66 instances, orange box in cropped and full confusion matrix in Fig. 5e and Fig. 4b). In this case we note that the bright spot corresponding to the hot engine panel at the back of the ZSU23 in the orientation $105^{\circ}-110^{\circ}$ is being confused with the hot rear exhaust of the BTR70 in the orientation $125^{\circ}-130^{\circ}$. To disrupt the prominence of these common features we make the hot and bright engine on the rear of the ZSU23 significantly darker, in contrast to the real samples, reducing the realism of the synthetic data. This disruptive modification results in an increase in mAP50 of 1.1% from 94.6% to 95.7% (Fig. 4b,d, Table 1), similar to that achieved from using the reinforcing modification above.
@@
Reinforcing + Disruptive Modifications: When we combine both modifications at the same time, we observe a reduction in the misclassification of BTR70 with both SUV and ZSU23 (Fig. 4e), and a corresponding increase in mAP50 of 1.5% on average from 94.6% to 96.1%, which is larger than the 1.1% increase in performance gained from either modification alone (Table 1).
@@ 6 Discussion
We have demonstrated that object detection models trained on real data, supplemented with synthetic data generated from 3D mesh models, can be improved using XAI-guided modifications to the mesh models themselves. Importantly, this framework allows both modifications that increase and decrease realism in synthetic data, through reinforcing unique features that distinguish between classes and through disrupting common features between classes. We show that both types of modifications can be used to reduce misclassifications and improve model performance. As a first proof of concept this demonstrates the power of XAI-driven modifications to improve synthetic datasets.
@@
In our current approach, considerable time, effort, and expertise is required to modify each mesh model. For future extension of this work, we therefore propose that the mesh model modifications could be automated by mapping the 2D SHAP images to the 3D mesh model, e.g. through use of ray-tracing to identify the mesh faces that constitute the unique and common features [59], and then setting the material properties of these faces using Probuilder [51]. The algorithm could scan through a large number of material property modifications on the selected mesh faces, and test the effect of each modification on the model performance. After training over a number of curated examples, an additional model could learn which modifications lead to the largest increases in performance, which would fully automate our data modification pipeline. Furthermore, while we currently use SHAP to show the location of common/unique features in the input image, mechanistic interpretability [41] could be used to also show the nature of these features (shape, texture, imperfections, etc.), which would allow us to more efficiently and effectively modify our mesh models.
@@
Introducing synthetic samples for unseen scenarios and steering a model’s attention in specific directions also presents a novel method to obtain more robust, unbiased, and safe models for critical applications such as autonomous driving [60] and medical diagnosis [61], which are sensitive to edge cases and class confusion. For example, an autonomous driving computer vision model could incorrectly associate a human walking pose with the basic task of detecting pedestrians, and therefore fail to detect stationary pedestrians on the road. Similarly, the same model could be biased towards particular racial, age or gender characteristics [60]. Therefore, interpreting the location and nature of such a model’s focus and procedurally generating appropriate synthetic samples to counter these effects could allow us to steer the model to instead focus on more relevant features for detecting pedestrians, and thus improve its robustness and generalisation.
@@
Steering model decisions through careful, explainable AI-guided curation of synthetic data enables us to identify and fix bugs, flaws and systematic biases in computer vision models, and is a path to achieving safer, robust and more responsible AI systems.
Acknowledgement: Research was sponsored by the Army and was accomplished under Cooperative Agreement Number W911NF-22-2-0161. The views and conclusions contained in this document are those of the authors and should not be interpreted as representing the official policies, either expressed or implied, of the U.S. Army or the U.S. Government. The U.S. Government is authorized to reproduce and distribute reprints for Government purposes notwithstanding any copyright notation herein.
@@ 9 Confusion Matrices
Figure 8: Confusion matrices for the model trained on only real data from the dataset (i.e no synthetic data). Panels (a)-(d) correspond to 4 different seeds and panel (e) corresponds to the average over these 4 different seeds and is the same confusion matrix as that which appears in panel a of Fig. 4a in the main paper.
@@
For each trained model in the main paper (reported in Table 1 ) we provided average confusion matrices (Fig. 4 in main paper). Here, in Supplementary Figs. 8-12, we provide the complete set of four confusion matrices that were used to calculate this average, corresponding to the four random seeds used for each trained model.
@@
Figure 9: Confusion matrices for the model trained on real data and the initial synthetic data (dataset v0). Panels (a)-(d) correspond to 4 different seeds and panel (e) corresponds to the average over these 4 different seeds and is the same confusion matrix as that which appears Fig. 4b in the main paper.
@@
Figure 10: Confusion matrices for the model trained on real data and synthetic data after the reinforcing modification (dataset vR) has been applied. Panels (a)-(d) correspond to 4 different seeds and panel (e) corresponds to the average over these 4 different seeds and is the same confusion matrix as that which appears in Fig. 4c in the main paper.
@@
Figure 11: Confusion matrices for the model trained on real data and synthetic data after the disruptive modification (dataset vD) has been applied. Panels (a)-(d) correspond to 4 different seeds and panel (e) corresponds to the average over these 4 different seeds and is the same confusion matrix as that which appears in Fig. 4d in the main paper.
@@
Figure 12: Confusion matrices for the model trained on real data and synthetic data after the reinforcing and disruptive modifications (dataset v(R+D)) have been applied. Panels (a)-(d) correspond to 4 different seeds and panel (e) corresponds to the average over these 4 different seeds and is the same confusion matrix as that which appears in Fig. 4e in the main paper.

### v2

@@ 4 Experimental Setup
Our aim is to perform object detection, that is, to detect and classify any vehicle(s) in the image, and estimate the bounding box coordinates of any detected vehicles. We use a YOLOv8 nano (You Only Look Once) model by Ultralytics [59] that has been pretrained on the COCO dataset, and fine-tune it on the ATR DSIAC dataset.
In order to prepare the dataset, we first crop the input images of initial size height 512, width 640 to a size of $256$ by $512$ pixels, with the centre of the crop positioned at a random pixel in the input image, which also provides negative samples where no vehicle is present. The cropped image is then resampled using bilinear interpolation to a size of $128$ by $256$ pixels. We split the images according to vehicle orientations. For our test and real training datasets, we use ATR DSIAC images with vehicle orientations in the ranges $\mathcal{O}_{test}=[70^{\circ},110^{\circ}]\cup[250^{\circ},290^{\circ}]$ and $\mathcal{O}_{train}=[-20^{\circ},20^{\circ}]\cup[160^{\circ},200^{\circ}]$ respectively. For synthetic training data, vehicle orientations are in the range $\mathcal{O}_{syn}=[50^{\circ},130^{\circ}]\cup[230^{\circ},310^{\circ}]$, as illustrated in Fig. 3. We employ this train-test split so that we can examine how the addition and modification of synthetic data improve our model’s ability to detect orientations of target vehicles that have not been seen in training; synthetic data have especial utility for detection of unseen scenarios. This split also reduces data leakage between the train and test data, which can occur if train and test frames are sampled uniformly from the video dataset, as per much of the previous literature on the ATR DSIAC dataset [38, 29, 49, 35, 6]. As neighbouring frames in a video are highly correlated, results reported on such uniformly sampled train-test splits can be unrealistically high due to overfitting.
@@ 5 Results
To provide a baseline reference for improvements in performance we first fine-tune our YOLOv8n model using only the real data from the dataset, trained using $9000$ images from $\mathcal{O}_{train}$ and tested using $\mathcal{O}_{test}$, where there is no overlap between the vehicle orientations in $\mathcal{O}_{train}$ and $\mathcal{O}_{test}$. We report performance using the mean average precision score at an intersection over union (IoU) threshold of $0.5$ (mAP50). We take ensemble averages over 4 different random seeds to demonstrate that our model improvements are robust to seed variations. The baseline performance for YOLOv8n when the model is only trained on the real dataset shows an mAP50 score of $90.0\%$ (Table 1, first row). This increases to $94.6\%$ when we instead train the YOLOv8n model on the same real data plus 9000 synthetic images generated from our initial mesh models, which we term v0 (version 0) and which includes vehicle orientations ($\mathcal{O}_{syn}$) that overlap the test time orientations in the real data $\mathcal{O}_{test}$ (compare Fig. 3 centre and right panels). Through a suite of tests we show that this increase in performance is not simply due to increased dataset size, and that the optimal ratio of real to synthetic data is 0.5 (supplementary Fig. 17), which we adopt through the rest of our experiments.
@@
YOLOv8n YOLOv8s YOLOv8x
Real 90.0% - -
Real + Syn v0 (Initial) 94.6% 95.6% 89.6%
Real + Syn v0 (Varifocal Loss) 94.5% - -
Real + Syn v0 (Online Hard Example Mining) 94.4% - -
Real + Syn vR (Reinforcing) 95.7% 96.4% 91.6%
Real + Syn vD (Disruptive) 95.7% 95.8% 89.9%
Real + Syn v(R+D) (Reinforcing and Disruptive) 96.1% 96.4% 90.7%
@@ 5.1 Using XAI to improve model performance
To understand the increase in mAP50 score from the addition of the v0 synthetic data to the training dataset, we compare the confusion matrices in Fig. 8a,b and see a large reduction in the background being misclassified as vehicles (e.g. dark pink box). By plotting the SHAP values of the pixels, averaged over $100$ random correctly classified test images, we see an increase in focus inside the bounding box of the target vehicle (average SHAP values increase from $39\%$ to $54\%$, see supplementary Fig. 2), which indicates the reduction in misclassification of the background is due to increased model focus on the vehicles relative to the background, when synthetic v0 data are added in training. However, with the addition of synthetic data v0, the confusion matrices also show a small increase in confusion (decrease in performance) between some vehicles in the dataset e.g. BTR70 misclassified as SUV or ZSU23 (light pink and orange boxes respectively in Fig. 8b and Fig. 5b,e). In the following sections we target these misclassifications using our new procedure from Section 2.1, demonstrating the use of both Reinforcing and Disruptive modifications to the synthetic data v0 in order to reduce model confusion and improve performance.
@@
Figure 4: Confusion matrices calculated by taking an average over $4$ different seeds for models trained with real data plus (a) no synthetic data (b) synthetic data v0, (c) synthetic data v(R+D) which includes both reinforcing modification, vR, to unique features of SUV relative to BTR70 and disruptive modification, vD, to common features of ZSU23 and BTR70.
@@
Figure 5: Illustration of the synthetic data modification process for reducing confusion. a,d: Ground-truth samples. b,e: Top shows confusion matrix for model trained on the initial synthetic data showing misclassifications highlighted in the confusion matrix. Bottom shows synthetic data samples (left), and SHAP contribution plots (right) for correct classifications capturing the unique features, and for misclassifications that capture the common features. c,f: Top shows confusion matrix for model trained on modified synthetic data, showing reduced misclassifications (see Supplementary Fig. 3 for full confusion matrices). Bottom shows samples of modified synthetic data (left), and SHAP contribution plots (right) for correct classifications illustrating an increased focus on unique features and reduced focus on common features. Pixels with a contribution score $<50\%$ (chosen to represent a trade-off between information and noise) of the highest score are masked in purple.
@@
Reinforcing Modification Example: Fig. 5a-c shows how we use Reinforcing Modification to reduce misclassification of the SUV class as a BTR70 (57 instances, light pink box in cropped and full confusion matrix in Fig. 5b and Fig. 8b). To identify the cause of confusion we group the validation images into $5^{\circ}$ segments of vehicle orientations, which are this dataset’s natural dimension of variation, and plot three average SHAP maps within each cluster; one each for the correct classification of the SUV and BTR70 and one for the misclassification of the SUV as a BTR70. Correct classifications of the SUV have increased focus towards the rear of the vehicle (Fig. 5b, top green box), whereas misclassifications of the SUV as a BTR70 do not focus in this same area (Fig. 5b, bottom green box) and instead have increased focus on the front bonnet of the SUV, which is a brighter (hotter) area in the IR images. This bright area is in a similar location to the hot rear exhaust of the BTR70 when the SUV is orientated $70^{\circ}-75^{\circ}$ (see Supplementary Fig. 18a,b) and the BTR70 is orientated $260^{\circ}-265^{\circ}$ (Fig. 5b, red boxes, Supplementary Fig. 18c,d). The bright SUV bonnet and BTR70 exhaust therefore represent common features and the side and rear of the SUV represent unique features. We reinforce model focus on the unique features of the SUV by making the main frame and rear part of the synthetic SUV images more similar to the real SUV images, by reducing the smoothness parameter of the material on the SUV mesh model’s surface. This reduces reflections off the smooth SUV surface (visible as bright-to-dark gradients in Fig. 5b, blue box) and results in more uniform texture (Fig. 5c, blue box), matching the real image (Fig. 5a, top). The effect of this modification is to increase the focus on the unique features and accordingly reduce the focus on common features (Fig. 5c). This results in a reduction of SUV, BTR70 misclassifications from 57 to 24 and an mAP50 increase of 1.1% from 94.6% to 95.7% (Fig. 5b,c, Table 1).
@@
Disruptive Modification Example: Fig. 5d-f shows how we use Disruptive Modification to reduce misclassification of the ZSU23 class as a BTR70 (66 instances, see Fig. 8b,c). In this case we note that the bright spot corresponding to the hot engine panel at the back of the ZSU23 in the orientation $105^{\circ}-110^{\circ}$ (see Supplementary Fig. 18a,b) is being confused with the hot rear exhaust of the BTR70 in the orientation $110^{\circ}-120^{\circ}$ (see Supplementary Fig. 18c,d). To disrupt the prominence of these common features we make the hot and bright engine on the rear of the ZSU23 significantly darker, in contrast to the real samples, reducing the realism of the synthetic data. This disruptive modification results in an increase in mAP50 of 1.1% from 94.6% to 95.7% (Table 1), similar to that achieved from using the reinforcing modification above. We also note that both types of modification result in better mAP50 scores than from use of varifocal loss [66] or online hard example mining (OHEM, [53]) instead when training on the Real+Syn v0 dataset. These methods aim to improve classification performance of “hard examples” through optimising the training procedure, but these methods, used here as benchmarks, actually show a slight performance drop of 0.1-0.2% compared to baseline results for the model trained on the Real+Syn v0 dataset (Table 1, rows 2-6).
@@
Reinforcing + Disruptive: When we combine both modifications, we observe reduction in the misclassification of BTR70 with both SUV and ZSU23 (Fig. 8e), and a corresponding increase in mAP50 of 1.5% on average from 94.6% to 96.1%; larger than the 1.1% increase in performance gained from either modification alone (Table 1).
These results are similar when we repeat our experiments with YOLOv8s and YOLOv8x architectures; the same modifications to the training data all lead to improved mAP50 scores averaged across 4 random seeds (Table 1, final two columns). For these larger models, the reinforcing modification results in a larger mAP50 increase than the disruptive modification, but some variation is expected as for simplicity we have used the same modifications to the synthetic data that we derived from the YOLOv8n model - in reality we expect the exact nature of modifications required to vary with model architecture.
@@ 6 Discussion
We have demonstrated that XAI techniques can be used to guide modification of synthetic datasets used to train object detection models, in order to improve model performance. Importantly, our methodology reduces the workload of human operators in designing and optimising such datasets, and also allows both modifications that increase and decrease realism in synthetic data, through reinforcing unique features that distinguish between classes and through disrupting common features between classes. We show that both types of modifications can be used to reduce misclassifications and improve model performance. As a first proof of concept this demonstrates the power of XAI-driven modifications to generate better synthetic data more efficiently.
@@
Our current results already have potential to enable faster generation of synthetic datasets in a human-in-the-loop process. But although such human-in-the-loop processes are often desirable for maintaining control over trained deep learning models, future extension of this work could move towards partial automation. This could be through automated modification of the 3D mesh model by mapping the 2D SHAP images to the 3D mesh using ray-tracing [30], modifying the mesh corresponding to unique or common features by changing the material properties adding or removing sub-parts of the vehicle, rotating or scaling sub-parts, and then training an additional model to learn which modifications lead to the largest increases in performance. Alternatively our method is not limited to use of game engines, and could also work with diffusion models if they can provide the level of fine-grained control needed for modifications. Furthermore, while we currently use SHAP to show the location of common and unique features in the input image, mechanistic interpretability [4] could be used to also show the nature of these features (shape, texture, imperfections, etc.), which would allow us to more efficiently and effectively modify our mesh models.
@@
Introducing synthetic samples for unseen scenarios and steering a model’s attention in specific directions also presents a novel method to obtain more robust, unbiased, and safe models for critical applications such as autonomous driving [26] and medical diagnosis [32], which are sensitive to edge cases and class confusion, and have shown bias towards racial, age or gender characteristics [26]. Interpreting the location and nature of such a model’s focus and procedurally generating appropriate synthetic samples to counter these effects could allow us to steer the model to instead focus on more relevant features, and thus improve its robustness and generalisation.
@@
Steering model decisions through careful, explainable AI-guided curation of synthetic data enables us to identify and fix bugs, flaws and systematic biases in computer vision models, whilst simultaneously overcoming a challenge of critical practical importance; reducing the workload of human operators in designing and optimising synthetic datasets.
@@ 9 Confusion Matrices
In Section 5 in the main paper, we present results of the ATR dataset trained on different synthetic data modifications. The full confusion matrices of the initial confusion when the model is trained on real data only and the real data and synthetic data constructed from the initial mesh models is reproduced from Fig. 4a,b in the main text in Fig. 8a,b respectively. The remaining panels, c,d show the full confusion matrices corresponding to the truncated confusion matrices in Fig. 5c,f from the main paper for the reinforcing (vR) and disruptive (vD) modifications respectively. Again Fig. 8e is a reproduction of Fig. 4c from the main paper corresponding to the confusion matrix for the reinforcing and disruptive v(R+D) modifications.
Figure 8: Confusion matrices calculated by taking an average over $4$ different seeds for YOLOv8n models trained with real data plus (a) no synthetic data (b) synthetic data v0, (c) synthetic data vR (reinforcing modification to unique features of SUV relative to BTR70), (d) synthetic data vD (disruptive modification to common features of ZSU23 and BTR70), and (e) synthetic data v(R+D)
@@
Figure 9: Confusion matrices for the YOLOv8n model trained on only real data from the dataset (i.e no synthetic data). Panels (a)-(d) correspond to 4 different seeds and panel (e) corresponds to the average over these 4 different seeds and is the same confusion matrix as that which appears in panel a of Fig. 4a in the main paper.
@@
For each trained model in the main paper (reported in Table 1) we provided average confusion matrices (Fig. 4 in main paper). Here, in Supplementary Figs. 9-13, we provide the complete set of four confusion matrices that were used to calculate this average, corresponding to the four random seeds used for each trained model. We also provide the confusion matrices for different seeds for experiments with YOLOv8s and YOLOv8x in Fig. 14-20.
@@
Figure 10: Confusion matrices for the YOLOv8n model trained on real data and the initial synthetic data (dataset v0). Panels (a)-(d) correspond to 4 different seeds and panel (e) corresponds to the average over these 4 different seeds and is the same confusion matrix as that which appears Fig. 4b in the main paper.
@@
Figure 11: Confusion matrices for the YOLOv8n model trained on real data and synthetic data after the reinforcing modification (dataset vR) has been applied. Panels (a)-(d) correspond to 4 different seeds and panel (e) corresponds to the average over these 4 different seeds and is the same confusion matrix as that which appears in Fig. 4c in the main paper.
@@
Figure 12: Confusion matrices for the YOLOv8n model trained on real data and synthetic data after the disruptive modification (dataset vD) has been applied. Panels (a)-(d) correspond to 4 different seeds and panel (e) corresponds to the average over these 4 different seeds and is the same confusion matrix as that which appears in Fig. 4d in the main paper.
@@
Figure 13: Confusion matrices for the YOLOv8n model trained on real data and synthetic data after the reinforcing and disruptive modifications (dataset v(R+D)) have been applied. Panels (a)-(d) correspond to 4 different seeds and panel (e) corresponds to the average over these 4 different seeds and is the same confusion matrix as that which appears in Fig. 4e in the main paper.
Figure 14: Confusion matrices for the YOLOv8s model trained on real data and initial synthetic data (dataset v0). Panels (a)-(d) correspond to 4 different seeds and panel (e) corresponds to the average over these 4 different seeds.
Figure 15: Confusion matrices for the YOLOv8s model trained on real data and synthetic data after reinforcing modification (dataset vR) has been applied. Panels (a)-(d) correspond to 4 different seeds and panel (e) corresponds to the average over these 4 different seeds.
Figure 16: Confusion matrices for the YOLOv8s model trained on real data and synthetic data after disruptive modification (dataset vD) has been applied. Panels (a)-(d) correspond to 4 different seeds and panel (e) corresponds to the average over these 4 different seeds.
Figure 17: Confusion matrices for the YOLOv8s model trained on real data and synthetic data after reinforcing + disruptive modifications (dataset v(R+D)) have been applied. Panels (a)-(d) correspond to 4 different seeds and panel (e) corresponds to the average over these 4 different seeds.
Figure 18: Confusion matrices for the YOLOv8x model trained on real data and initial synthetic data (dataset v0). Panels (a)-(c) correspond to 3 different seeds and panel (d) corresponds to the average over these 3 different seeds.
Figure 19: Confusion matrices for the YOLOv8x model trained on real data and synthetic data after reinforcing modification (dataset vR) has been applied. Panels (a)-(c) correspond to 3 different seeds and panel (d) corresponds to the average over these 3 different seeds.
Figure 20: Confusion matrices for the YOLOv8x model trained on real data and synthetic data after disruptive modification (dataset vD) has been applied. Panels (a)-(c) correspond to 3 different seeds and panel (d) corresponds to the average over these 3 different seeds.
Figure 21: Confusion matrices for the YOLOv8x model trained on real data and synthetic data after reinforcing + disruptive modifications (dataset v(R+D)) have been applied. Panels (a)-(c) correspond to 3 different seeds and panel (d) corresponds to the average over these 3 different seeds.